Which of the following results in a recognized gain or loss?

- Kelly sells her vacation cabin (adjusted basis $100,000) forr $150,000.

- Adam sells his personal residence (adjusted basis $150,000) for $100,000.

-Carl's personal residence (adjusted basis $65,000) is condemned by the city. He receives condemnation proceeds of $55,000.

- Olga's land is worth $40,000 at the end of the year. She had purchased the land six months earlier for $25,000.

- Vera's personal car (adjusted basis $22,000) is stolen/ She receives $23,000 from the insurance company and does not plan to replace the car.

- Jerry sells used clothing (adjusted basis $500) to a thrift store for $50.
Answer #1

a. Recognized Gain

Realized Gain = Amount realized - Adjusted basis

= 150,000-100,000

= 50,000

b.Un recognized loss

Realized loss = 100,000-150,000

=(50,000)

c. Un recognized Loss

Realized Loss = Amount realized -Adjusted basis

   =55,00-65,00

=(10,000)

d. No loss or Gain (since no sale or disposition is made)

e.Recognized Gain

Realized Gain = Amount realized - Adjusted basis

   =23,000-22,000

=1,000

f. Un recognized loss

Realized loss = Amount realized -Adjusted basis

=50-500

   =(450)